A glass of relative permitivity 4 is kept in an external electric field of magnitude 10^2 Vm^–1
.
Calculate the polarisation vector, molecular/atomic polarisability and the refractive index
of the glass
1
Expert's answer
2020-02-20T09:31:58-0500

Polarization vector


"P=\chi \epsilon_0 E"


"\epsilon=1+\chi \to \chi=\epsilon-1=4-1=3"


"P=\chi \epsilon_0 E=3\cdot 8.85\cdot 10^{-12}\cdot 100=26.55\cdot 10^{-10} \frac{F\cdot V}{m^2}"


Molecular/atomic polarisability



"P=N\alpha \epsilon_0 E \to \alpha=\frac{P}{N \epsilon_0 E}=\frac{26.55\cdot 10^{-10}}{6.02\cdot 10^{23}\cdot 8.85\cdot 10^{-12}\cdot 100}=0.5\cdot 10^{-23} m^3"



Refractive index


"n=\sqrt{\epsilon\mu}=\sqrt{4\cdot 0.99}=1.99"